# On the Complexity of Minimizing Energy Consumption of Partitioning DAG Tasks

### v1

@@ 4 Energy-Saving Partition of DAG
$\begin{split}{\textsf{Tran}_{G^{\prime}}(f,q)}&=\sum_{i\in[k]}\sum_{x\in f^{-1}(i)}\sum_{j\in[k]\setminus\{i\}}q(x)\cdot{\bf{1}}_{G^{\prime}}(x,f^{-1}(j))\\ &=\sum_{i\in[k]}\sum_{v\in V_{i}}\sum_{\begin{subarray}{c}a(v,u)\in B(v),\\ j\in[k]\setminus\{i\},u\in f^{-1}(j)\end{subarray}}q(a(v,u))\\ &=\sum_{i\in[k]}\sum_{v\in V_{i}}\sum_{\begin{subarray}{c}(v,u)\in A(\overrightarrow{G}),\\ j\in[k]\setminus\{i\},u\in V_{j}\end{subarray}}q(a(v,u))\\ &=\sum_{\begin{subarray}{c}i,j\in[k],i\neq j,\\ v\in V_{i},u\in V_{j},\\ \{v,u\}\in E(G)\end{subarray}}w(\{v,u\})\leq r.\\ \end{split}$ (4)
To see that Equality (4) holds, observe that for every $v\in V_{i}$, where $i\in[k]$, it holds that $B(v)\subseteq B[V_{i}]$ and $\Gamma^{+}_{G^{\prime}}{(v)}=B(v)$. Consequently, only outputs of newly introduced vertices $a(v,u)\in V(G^{\prime})\setminus V(\overrightarrow{G})$ necessitate transfer. Moreover, the outdegree of each newly introduced vertex $a(v,u)$ is exactly one. This ensures the correctness of the transition from the first line to the third line in Equality (4). The transition from the penultimate line to the last line in Equality (4) follows from the construction of $\overrightarrow{G}$ from $G$, and the definition of the function $q$. Now, we can conclude that the instance of ESP-DAG is a Yes-instance.
()($\Leftarrow$)( ⇐ ) Assume that the instance of ESP-DAG is a Yes-instance, i.e., there is an assignment function $f:V(G^{\prime})\rightarrow[k]$ such that
$\left(\sum_{x\in V(G^{\prime})}p(x,f(x))\right)+{\textsf{Tran}_{G^{\prime}}(f,q)}\leq r.$ (5)
By Equality (3), for every terminal $t_{i}\in T$, where $i\in[k]$, it holds that $f(t_{i})=i$. As $p(x,i)=0$ for all $x\in V(G^{\prime})\setminus T$ and $i\in[k]$, we have that $\sum_{x\in V(G^{\prime})}p(x,f(x))=0$. For each $i\in[k]$, let $V_{i}=\{v\in V(G):f(v)=i\}$. We show below that the total weight of edges in $G$ crossing the partition $(V_{i})_{i\in[k]}$ is exactly ${\textsf{Tran}_{G^{\prime}}(f,q)}$, which is at most $r$. To this end, recall that $q(v)=+\infty$ for every $v\in V(G)$. Then, by Inequality (5), we know that $f(v)=f(x)$ for all $x\in B(v)$ and $v\in V(G)$. That is, for every $v\in V(G)$$v$ and all its outneighbors in $G^{\prime}$ are assigned the same value by $f$. To be more precise, for every $i\in[k]$, it holds that $B[V_{i}]\subseteq f^{-1}(i)$. As $B[V_{i}]$ and $B[V_{j}]$ are disjoint as long as $V_{i}$ and $V_{j}$ are, this indeed means that $B[V_{i}]=f^{-1}(i)$ for all $i\in[k]$. Then, by the same reasoning in the proof for the ()($\Rightarrow$)( ⇒ ) direction, we infer that Equality (4) holds too in this direction. Therefore, the instance of Multiway Cut is a Yes-instance. ∎
@@
ESP-DAG is NP-hard for every $k\geq 3$ even when the two functions $p$ and $q$ have overall two different values.
The proof of the result can be done by replacing the value 00 in Equality (3) with $1$, resetting $q(a(v,u))=1$ for all newly introduced vertices $a(v,u)$, and resetting $r\coloneqq r+|V|$ in the instance of ESP-DAG constructed in our reduction.
When $p$ is a constant function, ESP-DAG can be solved trivially. Therefore, we have a complexity dichotomy for ESP-DAG with respect to the number of different values of the function $p$. Towards a complexity dichotomy for ESP-DAG concerning the number of machines, we have the following result.
@@
Let $(G,p,q,r)$ be an instance of ESP-DAG where $G$ is a DAG, and $p:V(G)\times[2]\rightarrow\mathbb{R}_{\geq 0}$ and $q:V(G)\rightarrow\mathbb{R}_{\geq 0}$ are two functions. Our reduction is as follows. We first construct a digraph $G^{\prime}$ obtained from $G$ by performing the following operations:
@@
      remove all arcs from $v$ to all its outneighbors except $(v,v^{+})$.
@@
    Create two vertices $s$ and $t$, add arcs from $s$ to all vertices in $V(G)$, and add arcs from all vertices in $V(G)$ to $t$. For each $v\in V(G)$, let the weight of the arc $(s,v)$ be $p(v,2)$, and let that of $(v,t)$ be $p(v,1)$. (Note that there are no arcs between the newly created vertices and $\{s,t\}$.)
Let $w:A(G^{\prime})\rightarrow\mathbb{R}_{\geq 0}$ be the function such that for each $e\in A(G^{\prime})$ it holds that $w(e)$ equals the weight of the arc $e$ defined above. The instance of DM-$s$-$t$-Cut is $(G^{\prime},w,s,t,r)$. The reduction clearly can be carried out in time $O(n+m)$, where $n$ and $m$ are respectively the number of vertices and the number of edges of $G$. It remains to prove its correctness.
@@
  • 1.
    For each $i\in[2]$, let $V_{i}^{i}=\{v\in V(G):\Gamma^{+}_{G}{[v]}\subseteq V_{i}\}$ be the set of vertices $v\in V(G)$ such that $v$ and all its outneighbors in $G$ are in the same set $V_{i}$.
  • 2.
    For each $i\in[2]$, let $V_{i}^{i\star}=\{v^{+}:v\in V_{i}^{i},\Gamma^{+}_{G}{(v)}\neq\emptyset\}\cup\{v^{-}:v\in V_{i}^{i},\Gamma^{+}_{G}{(v)}\neq\emptyset\}$ be the set of vertices constructed for vertices from $V_{i}^{i}$ by Operation (1a).
  • 3.
    For each $i,j\in[2]$ such that $i\neq j$, let $V_{i}^{j}=\{v\in V_{i}:\Gamma^{+}_{G}{(v)}\cap V_{j}\neq\emptyset\}$ be the set of vertices from $V_{i}$ having at least one outneighbor from $V_{j}$ in $G$.
  • 4.
@@
The above defined ten sets are pairwise disjoint, and their union is exactly $V(G^{\prime})\setminus\{s,t\}$. Let
@@
It is not difficult to verify that $X\cap Y=\emptyset$ and $X\cup Y=V(G^{\prime})$. To conclude this part of the proof, we need to demonstrate that the total weight of edges crossing the bipartition $(X,Y)$ in $G^{\prime}$ is at most $r$. To achieve this, we identify the edges crossing this bipartition, relying on Claims 1 through 5 presented below.
@@
By the construction of $G^{\prime}$ and the definitions of $V_{1}^{1}$ and $V_{1}^{1\star}$, all outneighbors of vertices of $V_{1}^{1}$ in $G^{\prime}$ are contained in $V_{1}^{1\star}\cup\{t\}$. As $V_{1}^{1\star}\subseteq X$, none of $V_{1}^{1}$ has any outneighbors from $Y\setminus\{t\}$ in $G^{\prime}$.
For a vertex $v^{-}\in V_{1}^{1\star}$ where $v\in V_{1}^{1}$, the only outneighbor of $v^{-}$ in $G^{\prime}$ is $v$ which is clearly not in $Y$. For a vertex $v^{+}\in V_{1}^{1\star}$ where $v\in V_{1}^{1}$, the outneighbors of $v^{+}$ in $G^{\prime}$ are exactly the outneighbors of $v$ in the graph $G$ which are all contained in $V_{1}^{1}$. As $V_{1}^{1}\subseteq X$, none of the outneighbors of $v^{+}$ is from $Y$.
By the definition of $V_{1}^{2-}$, every $v^{-}\in V_{1}^{2-}$ has exactly one outneighbor $v$ in $G^{\prime}$ which is from $V_{1}^{2}$. Clearly, $V_{1}^{2}$ and $Y$ are disjoint. ∎
@@
By the construction of $G^{\prime}$, every $v\in V(G)$ has exactly two outneighbors $v^{+}$ and $t$ in $G^{\prime}$. By the definition of $V_{1}^{2}$, for every $v\in V_{1}^{2}$$v^{+}$ is contained in $V_{1}^{2+}$. ∎
@@
Let $v$ be any arbitrary vertex in $V_{2}^{2}$, if there are any. We consider first the case where $v$ is neither a sink nor a source of $G$. By the construction of $G^{\prime}$, it holds that $\Gamma^{-}_{G^{\prime}}{(v)}\setminus\{s\}=\{v^{-}\}\cup\{u^{+}:u\in\Gamma^{-}_{G}{(v)}\}$ (Operations (1c) and (1d)). Clearly, $v^{-}\in V_{2}^{2\star}$. Let $u$ be a vertex in $\Gamma^{-}_{G}{(v)}$. If $u\in X$, then as $u\in V_{1}^{2}$ and $v\in Y$, we know that $u\in V_{1}^{2}$. It follows that $u^{+}\in V_{1}^{2+}$. If $u\in Y$, then either $u^{+}\in V_{2}^{2\star}$ (when $u\in V_{2}^{2}$) or $u^{+}\in V_{2}^{1+}$ (when $u\in V_{2}^{1}$). In both cases, we have that $u^{+}\in Y$. Therefore, none of the inneighbors of $v$ in $G^{\prime}$ are from $X\setminus\{s\}$. In the case where $v$ is a sink but not a source of $G$, we have that $\Gamma^{-}_{G^{\prime}}{(v)}\setminus\{s\}=\{u^{+}:u\in\Gamma^{-}_{G}{(v)}\}$, in the case where $v$ is a source but not a sink of $G$, we have that $\Gamma^{-}_{G^{\prime}}{(v)}\setminus\{s\}=\{v^{-}\}$, and in the case where $v$ is an isolated vertex in $G$, we have that $\Gamma^{-}_{G^{\prime}}{(v)}\setminus\{s\}=\emptyset$. The proof for that none of the inneighbors of $v$ in $G^{\prime}$ is from $X\setminus\{s\}$ in these cases are the same as above.
Consider now a vertex $v^{+}\in V_{2}^{2\star}$ where $v\in V_{2}^{2}$. By the construction of $G^{\prime}$, we know that $v$ is the only inneighbor of $v^{+}$ in $G^{\prime}$. For a vertex $v^{-}\in V_{2}^{2\star}$ where $v\in V_{2}^{2}$, by the construction of $G^{\prime}$, we know that $\Gamma^{-}_{G^{\prime}}{(v^{-})}=\Gamma^{+}_{G}{(v)}$. As $v\in V_{2}^{2}$, we know that $\Gamma^{+}_{G}{(v)}\subseteq V_{2}$. Clearly, $V_{2}$ is the union of $V_{2}^{1}$ and $V_{2}^{2}$ which are both contained in $Y$. We can conclude now that none of the inneighbors of any vertex in $V_{2}^{2\star}$ is from $X\setminus\{s\}$.
Now we prove the claim for $V_{2}^{1+}$. Let $v^{+}$ be a vertex from $V_{2}^{1+}$. Obviously, $v\in V_{2}^{1}$. By the construction of $G^{\prime}$$v$ is the only inneighbor of $v^{+}$. As $V_{2}^{1}\subseteq Y$, the claim holds. ∎
@@
  • 1.
    ${{\textsf{cut}}}_{G^{\prime}}(\{s\},Y)$, i.e., the set of all arcs from $s$ to all vertices of $Y$ in the graph $G^{\prime}$.
  • 2.
    ${{\textsf{cut}}}_{G^{\prime}}(X,\{t\})$, i.e., the set of all arcs from all vertices in $X$ to $t$ in the graph $G^{\prime}$.
  • 3.
    ${{\textsf{cut}}}_{G^{\prime}}(V_{1}^{2},V_{1}^{2+})$, i.e., the set of arcs from $V_{1}^{2}$ to $V_{1}^{2+}$ in the graph $G^{\prime}$.
  • 4.
    ${{\textsf{cut}}}_{G^{\prime}}(V_{2}^{1-},V_{2}^{1})$, i.e., the set of arcs from $V_{2}^{1-}$ to $V_{2}^{1}$ in the graph $G^{\prime}$.
Armed with Claim 5, we examine the size of the cut from $X$ to $Y$ in $G^{\prime}$. By Operation (2) in the construction of $G^{\prime}$, we know that there is an arc from $s$ to every vertex in $V_{2}^{1}\cup V_{2}^{2}$, and there is no arc from $s$ to any other vertices in $Y\setminus(V_{2}^{1}\cup V_{2}^{2})$. Then, by Operation (1c) in the construction of $G^{\prime}$, we know that
@@
Now we analyze $w({{\textsf{cut}}}_{G^{\prime}}(V_{1}^{2},V_{1}^{2+}))$. As for every vertex $v^{+}\in V(G^{\prime})$ where $v\in V(G)$$v$ is the only inneighbor of $v^{+}$ in $G^{\prime}$, we have that
@@
By Claim 1, and the facts that $V_{1}=V_{1}^{1}\cup V_{1}^{2}$, $V_{2}\subseteq Y$, and that every vertex from $V_{1}^{2}$ has at least one outneighbor from $V_{2}$ in the graph $G$ (definition of $V_{1}^{2}$), we know that $\sum_{v\in V_{1}}q(v)\cdot{\bf{1}}_{G}(v,V_{2})=\sum_{v\in V_{1}^{2}}q(v)$ holds. Similarly, it holds that $\sum_{v\in V_{2}}q(v)\cdot{\bf{1}}_{G}(v,V_{1})=\sum_{v\in V_{2}^{1}}q(v)$. We arrive at
@@
$w({{\textsf{cut}}}_{G^{\prime}}(X,Y))=\sum_{v\in V_{i},i\in[2]}p(v,i)+{\textsf{BTran}_{G}((V_{1},V_{2}),q)}.$
Then, by Inequality (6), the instance of DM-$s$-$t$-Cut is a Yes-instance.
@@
Let $v\in V$ such that $\Gamma^{+}_{G}{(v)}\neq\emptyset$ and all outneighbors of $v$ in $G$ are in the same $Z\in\{X,Y\}$. Let $\overline{Z}=V(G^{\prime})\setminus Z$. Let $X^{\prime}=Z\cup\{v^{+},v^{-}\}$ and let $Y^{\prime}=\overline{Z}\setminus\{v^{+},v^{-}\}$. Then, it holds that $w({{\textsf{cut}}}_{G^{\prime}}(X^{\prime},Y^{\prime}))\leq w({{\textsf{cut}}}_{G^{\prime}}(X,Y))$ if $s\in X^{\prime}$, and $w({{\textsf{cut}}}_{G^{\prime}}(Y^{\prime},X^{\prime}))\leq w({{\textsf{cut}}}_{G^{\prime}}(X,Y))$ if $s\in Y^{\prime}$.

### v3

@@ 4 Energy-Saving Partition of DAG
$\begin{split}{\textsf{Tran}_{G^{\prime}}(f,q)}&=\sum_{i\in[k]}\sum_{v\in f^{-1}(i)}\sum_{j\in[k]\setminus\{i\}}q(v)\cdot{\bf{1}}_{G^{\prime}}(v,f^{-1}(j))\\ &=\sum_{i\in[k]}\sum_{v\in V_{i}}\sum_{\begin{subarray}{c}a(v,u)\in B(v),\\ j\in[k]\setminus\{i\},u\in f^{-1}(j)\end{subarray}}q(a(v,u))\\ &=\sum_{i\in[k]}\sum_{v\in V_{i}}\sum_{\begin{subarray}{c}(v,u)\in A(\overrightarrow{G}),\\ j\in[k]\setminus\{i\},u\in V_{j}\end{subarray}}q(a(v,u))\\ &=\sum_{\begin{subarray}{c}i,j\in[k],i\neq j,\\ v\in V_{i},u\in V_{j},\\ \{v,u\}\in E(G)\end{subarray}}w(\{v,u\})\leq r.\\ \end{split}$ (4)
To verify that Equality (4) holds, observe that for every $v\in V_{i}$, where $i\in[k]$, we have $B(v)\subseteq B[V_{i}]$ and $N^{+}_{G^{\prime}}{(v)}=B(v)$. Consequently, only the outputs of the newly introduced vertices $a(v,u)\in V(G^{\prime})\setminus V(\overrightarrow{G})$ necessitate transfer. Additionally, each newly introduced vertex $a(v,u)$ has exactly one outneighbor $u$. This guarantees the correctness of the transition from the first line to the third line in Equality (4). The transition from the penultimate line to the last line in Equality (4) follows from the construction of $\overrightarrow{G}$ from $G$ and the definition of the function $q$. Now, we can conclude that the instance of the ESP-DAG problem is a Yes-instance.
()($\Leftarrow$)( ⇐ ) Assume that the instance of the ESP-DAG problem is a Yes-instance, i.e., there is an assignment function $f:V(G^{\prime})\rightarrow[k]$ such that
$\left(\sum_{v\in V(G^{\prime})}p(v,f(v))\right)+{\textsf{Tran}_{G^{\prime}}(f,q)}\leq r.$ (5)
By Equality (3), for every terminal $t_{i}\in T$, where $i\in[k]$, we have $f(t_{i})=i$. Since $p(v,i)=0$ for all $v\in V(G^{\prime})\setminus T$ and $i\in[k]$, it follows that $\sum_{v\in V(G^{\prime})}p(v,f(v))=0$. For each $i\in[k]$, let $V_{i}=\{v\in V(G):f(v)=i\}$. We show below that the total weight of edges in $G$ crossing the partition $(V_{i})_{i\in[k]}$ is exactly ${\textsf{Tran}_{G^{\prime}}(f,q)}$, which is at most $r$. To this end, recall that $q(v)=+\infty$ for every $v\in V(G)$. By Inequality (5), we know that $f(v)=f(v^{\prime})$ for all $v\in V(G)$ and $v^{\prime}\in B(v)$. That is, for every $v\in V(G)$$v$ and all its outneighbors in $G^{\prime}$ are assigned the same value by $f$. To be more precise, for every $i\in[k]$, it holds that $B[V_{i}]\subseteq f^{-1}(i)$. Since $B[V_{i}]$ and $B[V_{j}]$ are disjoint whenever $V_{i}$ and $V_{j}$ are disjoint, this indeed means that $B[V_{i}]=f^{-1}(i)$ for all $i\in[k]$. Then, by the same reasoning in the proof of the ()($\Rightarrow$)( ⇒ ) direction, we infer that Equality (4) holds in this direction as well. Therefore, the instance of the Multiway Cut problem is a Yes-instance. ∎
@@
ESP-DAG is NP-hard for every $k\geq 3$ even when the two functions $p$ and $q$ have overall two different values: $1$ and +$\infty$+ ∞.
The proof of the result can be done by replacing the value 00 in Equality (3) with $1$, resetting $q(a(v,u))=1$ for all newly introduced vertices $a(v,u)$, and resetting $r\coloneqq r+|V(G)|$ in the instance of ESP-DAG constructed in our reduction.
When $p$ is a constant function, ESP-DAG can be solved trivially. Therefore, we have a complexity dichotomy for ESP-DAG with respect to the number of different values of the function $p$. Towards a complexity dichotomy concerning the number of machines, we have the following result.
@@
We derive a linear-time reduction from ESP-DAG with $k=2$ to DM-$s$-$t$-Cut as follows. Let $(G,p,q,r)$ be an instance of ESP-DAG where $G$ is a DAG, and $p:V(G)\times[2]\rightarrow\mathbb{R}_{\geq 0}$ and $q:V(G)\rightarrow\mathbb{R}_{\geq 0}$ are two functions. We first construct a digraph $G^{\prime}$ obtained from $G$ by performing the following operations:
@@
      remove all arcs from $v$ to all its outneighbors in $G$. Therefore, $V(G)$ forms an independent set in $G^{\prime}$.
@@
    Create two vertices $s$ and $t$, add arcs from $s$ to all vertices in $V(G)$, and add arcs from all vertices in $V(G)$ to $t$. For each $v\in V(G)$, let the weight of the arc $(s,v)$ be $p(v,2)$, and let that of $(v,t)$ be $p(v,1)$.
We refer to Figure 3 for an illustration of the construction of $G^{\prime}$.
Figure 3: Construction of the digraph $G^{\prime}$ from the DAG $G$ as described in the proof of Theorem 2. Red arcs have positive infinite weights.
Let $w:A(G^{\prime})\rightarrow\mathbb{R}_{\geq 0}$ be the function such that for each arc $e\in A(G^{\prime})$ it holds that $w(e)$ equals the weight of $e$ defined above. The instance of DM-$s$-$t$-Cut is $(G^{\prime},w,\{s,t\},r)$. The reduction clearly can be carried out in linear time. It remains to prove its correctness.
@@
  • For each $i\in[2]$, let $V_{i}^{i}=\{v\in V(G):N^{+}_{G}{[v]}\subseteq V_{i}\}$ be the set of vertices $v\in V(G)$ such that $v$ and all its outneighbors in $G$ are in the same set $V_{i}$. (Notice that each isolated vertex of $G$ is either in $V_{1}^{1}$ or in $V_{2}^{2}$.)
  • For each $i\in[2]$, let $V_{i}^{i\star}=\{v^{+}:v\in V_{i}^{i},N^{+}_{G}{(v)}\neq\emptyset\}\cup\{v^{-}:v\in V_{i}^{i},N^{+}_{G}{(v)}\neq\emptyset\}$ be the set of vertices constructed for vertices from $V_{i}^{i}$ by Operation (1a).
  • For each $i,j\in[2]$ such that $i\neq j$, let $V_{i}^{j}=\{v\in V_{i}:N^{+}_{G}{(v)}\cap V_{j}\neq\emptyset\}$ be the set of vertices from $V_{i}$ having at least one outneighbor from $V_{j}$ in $G$.
@@
The above defined ten sets are pairwise disjoint, and their union is exactly $V(G^{\prime})\setminus\{s,t\}$. Noticeably, $V_{1}=V_{1}^{1}\cup V_{1}^{2}$ and $V_{2}=V_{2}^{2}\cup V_{2}^{1}$. Let
@@
It is not difficult to verify that $X\cap Y=\emptyset$ and $X\cup Y=V(G^{\prime})$. Moreover, it holds that $V_{1}\subseteq X$ and $V_{2}\subseteq Y$. To conclude this part of the proof, we need to demonstrate that the total weight of edges crossing the bipartition $(X,Y)$ in $G^{\prime}$ is at most $r$. To achieve this, we identify the edges crossing this bipartition, relying on Claims 15 presented below.
@@
[…]
Consider first the set $V_{1}^{1}$. Let $v$ be an arbitrary vertex from $V_{1}^{1}$, if any. Apart from $t$, the outneighbors of $v$ in $G^{\prime}$ consist of the newly introduced vertex $v^{+}$ for $v$ (if $v$ is not a sink in $G$), and the newly introduced vertices $u^{-}$ for inneighbors $u$ of $v$ in $G$ (if $v$ is not a source of $G$). Obviously, if $v^{+}$ exists, it belongs to the set $V_{1}^{1\star}$, which is disjoint from $Y$. Additionally, any $u^{-}$ where $(u,v)\in A(G)$ cannot belong to $V_{2}^{2}\cup V_{2}^{1}\cup V_{1}^{2+}\cup V_{2}^{1+}$. Furthermore, $u^{-}$ cannot be in $V_{2}^{2\star}$ because this would imply $v\in V_{2}$ contradicting the assumption that $v\in V_{1}^{1}\subseteq V_{1}$. Thus, $v$ has no outneighbor from $Y\setminus\{t\}$ in $G^{\prime}$. Since $v$ was chosen arbitrarily from $V_{1}^{1}$, we conclude that no vertex in $V_{1}^{1}$ has any outneighbors from $Y\setminus\{t\}$ in $G^{\prime}$.
Next, consider the vertices in the set $V_{1}^{1\star}$. For every vertex $v^{-}\in V_{1}^{1\star}$ where $v\in V_{1}^{1}$, the only outneighbor of $v^{-}$ in $G^{\prime}$ is $v$, which is clearly not in $Y$. For a vertex $v^{+}\in V_{1}^{1\star}$ where $v\in V_{1}^{1}$, the outneighbors of $v^{+}$ in $G^{\prime}$ are exactly the outneighbors of $v$ in the digraph $G$, all of which belong to $V_{1}^{1}$. As $V_{1}^{1}\subseteq X$, none of the outneighbors of $v^{+}$ is from $Y$.
Finally, consider the set $V_{1}^{2-}$. By definition, each vertex $v^{-}\in V_{1}^{2-}$ has exactly one outneighbor $v$ in $G^{\prime}$, and $v$ is from $V_{1}^{2}$. Clearly, $V_{1}^{2}$ and $Y$ are disjoint. ∎
@@
The claim is vacuously true if $V_{1}^{2}=\emptyset$. Otherwise, let $v$ be an arbitrary vertex from $V_{1}^{2}$. The outneighbors of $v$ in $G^{\prime}$ are from $\{v^{+},t\}\cup\{u^{-}:(u,v)\in A(G)\}$. By the definition of $V_{1}^{2}$, it follows that $v^{+}$, if exists (when $v$ is not a sink in $G$), is contained in $V_{1}^{2+}$. Now suppose $v$ has an outneighbor $u^{-}$ in $G^{\prime}$ such that $(u,v)\in A(G)$. In this case, $u^{-}$ cannot be from $Y\setminus V_{2}^{2\star}=V_{2}^{2}\cup V_{2}^{1}\cup V_{1}^{2+}\cup V_{2}^{1+}\cup\{t\}$. If $u^{-}\in V_{2}^{2\star}$, then $u$ must be from $V_{2}^{2}$, which implies that $v\in V_{2}$, contradicting the fact that $v\in V_{1}^{2}\subseteq V_{1}$. Therefore, $v$ can only have $v^{+}$ and $t$ as outneighbors in $G^{\prime}$. As $v^{+}\in V_{1}^{2+}$, the claim holds. ∎
@@
We will prove the claim for the sets $V_{2}^{2}$, $V_{2}^{2\star}$, and $V_{2}^{1+}$ separately.
Consider first the set $V_{2}^{2}$. Let $v$ be an arbitrary vertex in $V_{2}^{2}$, if any. We consider first the case where $v$ is neither a sink nor a source of $G$. By the construction of $G^{\prime}$, it holds that $N^{-}_{G^{\prime}}{(v)}\setminus\{s\}=\{v^{-}\}\cup\{u^{+}:u\in N^{-}_{G}{(v)}\}$. Clearly, $v^{-}\in V_{2}^{2\star}$. Let $u$ be a vertex in $N^{-}_{G}{(v)}$. If $u\in X$, then $u$ belongs to $V_{1}^{2}$ because $u$ has an outneighbor $v\in V_{2}^{2}\subseteq V_{2}$. It follows that $u^{+}\in V_{1}^{2+}$, which is contained in $Y$. If $u\in Y$, then either $u^{+}\in V_{2}^{2\star}$ (when $u\in V_{2}^{2}$) or $u^{+}\in V_{2}^{1+}$ (when $u\in V_{2}^{1}$). In both cases, we have that $u^{+}\in Y$. We can conclude now that none of the inneighbors of $v$ in $G^{\prime}$ is from $X\setminus\{s\}$. In the case where $v$ is a sink but not a source of $G$, we have $N^{-}_{G^{\prime}}{(v)}\setminus\{s\}=\{u^{+}:u\in N^{-}_{G}{(v)}\}$. If $v$ is a source but not a sink of $G$, then $N^{-}_{G^{\prime}}{(v)}\setminus\{s\}=\{v^{-}\}$. If $v$ is an isolated vertex in $G$, we have $N^{-}_{G^{\prime}}{(v)}\setminus\{s\}=\emptyset$. The proofs for these cases follow the same reasoning as above.
Consider now a vertex $v^{+}\in V_{2}^{2\star}$, where $v\in V_{2}^{2}$. By the construction of $G^{\prime}$, we know that $v$ is the only inneighbor of $v^{+}$ in $G^{\prime}$. For a vertex $v^{-}\in V_{2}^{2\star}$, where $v\in V_{2}^{2}$, we have $N^{-}_{G^{\prime}}{(v^{-})}=N^{+}_{G}{(v)}$. As $v\in V_{2}^{2}$, we have $N^{+}_{G}{(v)}\subseteq V_{2}=V_{2}^{1}\cup V_{2}^{2}$. Since $V_{2}^{1}\cup V_{2}^{2}$ and $X$ are disjoint, we can conclude that none of the inneighbors of any vertex from $V_{2}^{2\star}$ belong to $X\setminus\{s\}$.
Finally, we prove the claim for $V_{2}^{1+}$. Let $v^{+}$ be a vertex from $V_{2}^{1+}$. By the construction of $G^{\prime}$$v$ is the only inneighbor of $v^{+}$. Obviously, $v\in V_{2}^{1}$, which is disjoint from $X$. ∎
@@
  • ${{\textsf{cut}}}_{G^{\prime}}(\{s\},Y)$, i.e., the set of all arcs from $\{s\}$ to $Y$ in the digraph $G^{\prime}$.
  • ${{\textsf{cut}}}_{G^{\prime}}(X,\{t\})$, i.e., the set of all arcs from $X$ to $\{t\}$ in the digraph $G^{\prime}$.
  • ${{\textsf{cut}}}_{G^{\prime}}(V_{1}^{2},V_{1}^{2+})$, i.e., the set of all arcs from $V_{1}^{2}$ to $V_{1}^{2+}$ in the digraph $G^{\prime}$.
  • ${{\textsf{cut}}}_{G^{\prime}}(V_{2}^{1-},V_{2}^{1})$, i.e., the set of all arcs from $V_{2}^{1-}$ to $V_{2}^{1}$ in the digraph $G^{\prime}$.
By Operation (2) in the construction of $G^{\prime}$, there is an arc from $s$ to every vertex in $V_{2}^{1}\cup V_{2}^{2}$, and there is no arc from $s$ to any vertices from $Y\setminus(V_{2}^{1}\cup V_{2}^{2})$. That is, ${{\textsf{cut}}}_{G^{\prime}}(\{s\},Y)$ consists of exactly the arcs from $s$ to all vertices of $V_{2}$ in $G^{\prime}$. Then, by Operation (1c) in the construction of $G^{\prime}$, we have
@@
Now we analyze $w({{\textsf{cut}}}_{G^{\prime}}(V_{1}^{2},V_{1}^{2+}))$. As for every $v^{+}\in V(G^{\prime})$ where $v\in V(G)$$v$ is the only inneighbor of $v^{+}$ in $G^{\prime}$, we have that
@@
By Claim 1, and the facts that $V_{1}=V_{1}^{1}\cup V_{1}^{2}$, $V_{2}\subseteq Y$, and every vertex from $V_{1}^{2}$ has at least one outneighbor from $V_{2}$ in $G$, we have that $\sum_{v\in V_{1}}q(v)\cdot{\bf{1}}_{G}(v,V_{2})=\sum_{v\in V_{1}^{2}}q(v)$. Similarly, it holds that $\sum_{v\in V_{2}}q(v)\cdot{\bf{1}}_{G}(v,V_{1})=\sum_{v\in V_{2}^{1}}q(v)$. We arrive at
@@
$w({{\textsf{cut}}}_{G^{\prime}}(X,Y))=\left(\sum_{v\in V_{i},i\in[2]}p(v,i)\right)+{\textsf{BTran}_{G}((V_{1},V_{2}),q)},$
which is at most $r$ by Inequality (6). Therefore, the instance of the DM-$s$-$t$-Cut problem is a Yes-instance.
@@
The first property, as formally articulated in the ensuing claim, affirms that when a vertex $v$ in $G$ is included in the same part of $(X,Y)$ alongside all its outneighbors in $G$, retaining both newly introduced vertices for $v$ within the same part as $v$ emerges as an optimal strategy.
@@
Let $v\in V(G)$ such that $N^{+}_{G}{(v)}\neq\emptyset$ and all outneighbors of $v$ in $G$ are in the same $Z\in\{X,Y\}$. Let $\overline{Z}=V(G^{\prime})\setminus Z$. Let $X^{\prime}=Z\cup\{v^{+},v^{-}\}$ and let $Y^{\prime}=\overline{Z}\setminus\{v^{+},v^{-}\}$. Then, it holds that $w({{\textsf{cut}}}_{G^{\prime}}(X^{\prime},Y^{\prime}))\leq w({{\textsf{cut}}}_{G^{\prime}}(X,Y))$ if $s\in X^{\prime}$, and $w({{\textsf{cut}}}_{G^{\prime}}(Y^{\prime},X^{\prime}))\leq w({{\textsf{cut}}}_{G^{\prime}}(X,Y))$ if $s\in Y^{\prime}$.